# Large Language Models Compression via Low-Rank Feature Distillation

### v1

@@ Abstract
Current LLM structured pruning methods involve two steps: (1) compressing with calibration data and (2) continued pretraining on billions of tokens to recover the lost performance. This costly second step is needed as the first step significantly impacts performance. Previous studies have found that pretrained Transformer weights aren’t inherently low-rank, unlike their activations, which may explain this performance drop. Based on this observation, we introduce a one-shot compression method that locally distills low-rank weights. We accelerate convergence by initializing the low-rank weights with SVD and using a joint loss that combines teacher and student activations. We reduce memory requirements by applying local gradient updates only. Our approach can compress Mixtral-8x7B within minutes on a single A100 GPU, removing 10 billion parameters while maintaining over 95% of the original performance. Phi-2 3B can be compressed by 40% using only 13 million calibration tokens into a small model that competes with recent models of similar size. We show our method generalizes well to non-transformer architectures: Mamba-3B can be compressed by 20% while maintaining 99% of its performance [^1] Code available at https://github.com/yaya-sy/SlimLlama.
@@
Large Language Models Compression via Low-Rank Feature Distillation
Figure 1: Our compression approach: STEP 1 selects layers to compress for a target compression ratio (e.g., N%) using various strategies (see Section 5). STEP 2 compresses and initializes the chosen parameters via SVD. STEP 3 distills the low-rank weights with a small calibration dataset.

### v2

# Lillama: Large Language Models Compression via Low-Rank Feature Distillation

@@ Abstract
Current LLM structured pruning methods typically involve two steps: (1) compression with calibration data and (2) costly continued pretraining on billions of tokens to recover lost performance. This second step is necessary as the first significantly impacts model accuracy. Prior research suggests pretrained Transformer weights aren’t inherently low-rank, unlike their activations, which may explain this drop. Based on this observation, we propose Lillama, a compression method that locally distills activations with low-rank weights. Using SVD for initialization and a joint loss combining teacher and student activations, we accelerate convergence and reduce memory use with local gradient updates. Lillama compresses Mixtral-8x7B within minutes on a single A100 GPU, removing 10 billion parameters while retaining over 95% of its original performance. Phi-2 3B can be compressed by 40% with just 13 million calibration tokens, resulting in a small model that competes with recent models of similar size. The method generalizes well to non-transformer architectures, compressing Mamba-3B by 20% while maintaining 99% performance [^1] Code available at https://github.com/yaya-sy/lillama.
@@
Lillama: Large Language Models Compression via Low-Rank Feature Distillation
Figure 1: Lillama approach: STEP 1 selects layers to compress for a target compression ratio (e.g., N%) using various strategies (see Section 5). STEP 2 compresses and initializes the chosen parameters via SVD. STEP 3 distills the low-rank weights with a small calibration dataset.